A linear regression model found the following :
Dependent variable : Quantity
Independent variables : X1 X2
coefficient
constant. 10
price. -2
Income. 3
R^2 = 0.83 t = 2.36
a. write the demand function as an equation
b. do the sign of the coefficients make sense ? why?
c. if price = 10, Income = 24 what is the predicted quantity sold?
d. find the point price elasticity at price =10, Income = 24
Answer #1

a) Demand Function:

Quantity demanded = 10 - 2*Price + 3*Income

b) Coefficient of the price is negative implying that price and quantity demanded are negatively related. Coefficient of Income is positive implying that income and quantity demanded are positively related.

c) Given, Price=10 and Income=24

Quantity demanded= 10 - (2*10) + (3*24)

= 10 - 20 + 72

=62 units